I need help with this problem
x(2x-3)=(2x+1)(x-4)

Found 2 solutions by scott8148, edjones:
Answer by scott8148(6628)

expanding both sides gives 2x^2-3x=2x^2-7x-4

subtracting 2x^2 gives -3x=-7x-4

adding 7x gives 4x=-4

Answer by edjones(8007)

x(2x-3)=(2x+1)(x-4)
2x^2-3x=2x^2-7x-4
4x=-4
x=-1
.
Ed